# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v25

@@ Abstract
Sinhala is the native language of the Sinhalese people who make up the largest ethnic group of Sri Lanka. The language belongs to the globe-spanning language tree, Indo-European. However, due to poverty in both linguistic and economic capital, Sinhala, in the perspective of Natural Language Processing tools and research, remains a resource-poor language which has neither the economic drive its cousin English has nor the sheer push of the law of numbers a language such as Chinese has. A number of research groups from Sri Lanka have noticed this dearth and the resultant dire need for proper tools and research for Sinhala natural language processing. However, due to various reasons, these attempts seem to lack coordination and awareness of each other. The objective of this paper is to fill that gap of a comprehensive literature survey of the publicly available Sinhala natural language tools and research so that the researchers working in this field can better utilize the contributions of their peers. As such, we shall be uploading this paper to arXiv and perpetually update it periodically as a living research article to reflect the advances made in the field. This manuscript is at version $5.7.0$.
@@
Sinhala, Natural Language Processing, Resource Poor Language
@@ 1 Introduction
Sinhala [^1] Englebretson and Genetti [1] observe that in some contexts the Sinhala language is also referred as Sinhalese, Singhala, and Singhalese language, being the native language of the Sinhalese people [2, 3, 4], who make up the largest ethnic group of the island country of Sri Lanka, enjoys being reported as the mother tongue (L1) of approximately 16 million people [5, 6]. When both L1 and L2 speakers are counted, $79.7\%$ of the total Sri Lankan population are literate in Sinhala [7]. A strong correlation between the island of Sri Lanka and the usage of Sinhala has been observed [8] [^2] Dunn and Edwards-Brown [8] in their work to augment Language Identification (LID) task with geography observe that for Sinhala LID does not benefit from the geography model given that the to models already has an alignment of 99%.. This implies two things: 1) the majority of the Sinhala linguistic sources are located in the geographical area of Sri Lanka, and 2) the majority of the available linguistic sources in Sri Lanka are in Sinhala.
@@
To give a brief linguistic background for the purpose of aligning the Sinhala language with the baseline of English, primarily it should be noted that the Sinhala language belongs to the same Indo-European language tree [9, 10, 11]. However, unlike English, which is part of the Germanic branch, Sinhala belongs to the Indo-Aryan branch. An extremely simplified version of the Indo-European language tree focusing mainly on Sinhala and its close relatives is shown in Fig 1. Further, Sinhala, unlike English, which borrowed the Latin alphabet, has its own writing system, which is a descendant of the Indian Brahmi script [12, 13, 14, 15, 16, 17, 18]. By extension, this makes Sinhala Script a member of the Aramaic family of scripts [19, 20]. Thus by inheritance, the Sinhala writing system is abugida (alphasyllabary), which to say, that consonant-vowel sequences are written as single units [21]. An extremely simplified evolution tree of the Sinhala script is shown in Fig 2. In it, the Sinhala script is shown as a singular derivation from Brahmi directly appearing with the contemporary script. But this is a simplification. We show the actual evolutionary eras of the Sinhala script in Appendix A where the gradual change from Brahmi to the contemporary script is clearer. It should be noted that the modern Sinhala language has loanwords from languages such as Tamil, English, Portuguese, and Dutch due to various historical reasons [22]. An analysis of the distribution of Sinhala letters towards the beginning, middle, or end of words conducted in the style of proposed by ProoffReader [23] using the SinMin Sinhala corpus created by Upeksha et al. [24, 25] is given in Appendix B.
@@
Figure 1: A portion of the Indo-European language tree with a focus on Sinhala and its close relatives. Only the Elu sub-tree is shown exhaustively. All other sub-trees only list some languages for representation purposes chosen randomly or otherwise. Some of the languages shown as Extinct have archaeological proof (eg: Prakrit) while others were retroactively reconstructed by linguists by observing the commonalities of languages that they begot (eg: Proto-Indo-Aryan). The status of some languages has spread beyond linguistics to be a social and political question. In this depiction, we have attempted to side with academic consensus when available and the majority when not. As an example of one contentions point on this diagram note the positioning of Pali. While what we show here is the academically accepted relationship, there are other schools of thought: 1) Some argue that Pali is a constructed language (ConLang) of Sinhala and thus should be depicted as a child of Sinhala., 2) Others claim it to be a sibling of Sinhala, 3) Further some claim it to be either be the link between Elu and Sinhala or the parent of Elu, 4) Some claim it to be merged with Gandhari, and 5) Some claim it to be the same as Magadhi Prakrit given that in some Buddhist traditions, it is noted Buddha to have spoken Magadhi.
@@
Figure 2: A portion of the writing systems (scripts) tree with a focus on Sinhala and its close relatives. It is important to note that this is distinct from the language tree shown in Fig 1. The writing systems used by a language being related, not being related, or the degree to which they are related is not directly correlated to the relevant languages being related, not being related, or the degree to which they are related. As an example, note that the writing systems of Sinhala and Tamil are reasonably closely related. But Tamil is not shown on Fig 1 given that it does not belong to the Indo-European language family. It belongs to the Dravidian language family. A language that is shown as a single entity in Fig 1 may have multiple entries in this diagram. An example is the language shown as Hindustani splits into Hindi and Urdu where the former uses the Devanagari script and the latter a derivation of Arabic script. Conversely, languages that may be shown as distinct entities in the full language tree, may map to a singular entity in the script tree. An example is the fact that English which would go under Proto-Germanic and Italian which would go under Proto-Italic in and extended version of Fig 1, would both map to Latin in this diagram given that both of them use the Latin script.
@@
Natural Language Processing (NLP) is a broad area covering all computational processing and analysis of human languages. To achieve this end, NLP systems operate at different levels [35, 36, 37]. A graphical representation of NLP layers and application domains are shown in Figure 3. On one hand, according to Liddy [36], these systems can be categorized into the following layers; phonological, morphological, lexical, syntactic, semantic, discourse, and pragmatic. The phonological layer deals with the interpretation of language sounds. As such, it consists of mainly speech-to-text and text-to-speech systems. In cases where one is working with the written text of the language rather than speech, it is possible to replace this layer with tools which handle Optical Character Recognition (OCR) and language rendering standards (such as Unicode [38]). The morphological layer analyses words at their smallest units of meaning. As such, analysis on word lemmas and prefix-suffix-based inflection are handled in this layer. Lexical layer handles individual words. Therefore tasks such as Part of Speech (PoS) tagging happens here. The next layer, syntactic, takes place at the phrase and sentence level where grammatical structures are utilized to obtain meaning. Semantic layer attempts to derive the meanings from the word level to the sentence level. Starting with Named Entity Recognition (NER) at the word level and working its way up by identifying the contexts they are set in until arriving at overall meaning. The discourse layer handles meaning in textual units larger than a sentence. In this, the function of a particular sentence maybe contextualized within the document it is set in. Finally, the pragmatic layer handles contexts read into contents without having to be explicitly mentioned [35, 36]. Some forms of anaphora (coreference) resolution [39, 40, 41, 42, 43] fall into this application.
@@
The objective of this paper is to serve as a comprehensive survey on the state of natural language processing resources for the Sinhala language. The initial structure and content of this survey are heavily influenced by the preliminary surveys carried out by de Silva [28] and Wijeratne et al. [35]. However, our hope is to host this survey at arXiv as a perpetually evolving, living research article [45, 46] which continuously gets updated as new research and tools for Sinhala language are created and made publicly available. We also discuss how the non-compliance of policies to put data and code online [47], after the research is concluded and the paper is published, has negatively impacted the growth and sustainability of Sinhala NLP. Hence, it is our hope that this work will help future researchers who are engaged in Sinhala NLP research to conduct their literature surveys efficiently and comprehensively. For the success of this survey, we shall also consider the Sri Lankan NLP tools repository, lknlp [^4] https://github.com/lknlp/lknlp.github.io. This manuscript is at version $5.7.0$. The latest version of the manuscript can be obtained from arXiv [^5] https://arxiv.org/abs/1906.02358 or ResearchGate [^6] http://bit.ly/31AhvvR. Given that this work was initially published in 2019, when citing this work, it is recommended to cite the most recent version to avoid the reviewers questioning its continued contemporary relevance. An example of an updated reference is given in Code Block LABEL:lst:ref.
@article{de2025survey,
@@
journal={arXiv preprint arXiv:1906.02358v25},
year={2025}
@@
The remainder of this survey is organized as follows: Section 2 introduces some important properties and conventions of the Sinhala language, which are important for the development and understanding of Sinhala NLP. Section 3 discusses the various tools and research available for Sinhala NLP. In this section, we would discuss both pure Sinhala NLP tools and research as well as hybrid Sinhala-English work. We will also discuss research and tools which contributes to Sinhala NLP either along with or by the help of Tamil, the other official language of Sri Lanka. Section 4 gives a brief introduction to the primary language sources used by the studies discussed in this work. Finally, Section 5, concludes the survey.
@@ 2 Properties of the Sinhala Language
Before moving on to discussing Sinhala NLP resources, we shall give a brief introduction to some of the important properties of Sinhala language, which impact the development of Sinhala NLP resources. Sinhala grammar has two forms: written (literary) and spoken. These forms differ from each-other in their core grammatical structures [49, 1, 50]. The written form strictly adheres to the SOV (Subject, Object, and Verb) configuration [51, 52]. Further, in the written form, subject-verb agreement is enforced [53] such that, in order to be grammatically correct, the subject and the verb must agree in terms of: gender (male/female), number (singular/plural) and person (1st/2nd/3rd). However, in spoken Sinhala, the SOV order can be neglected [54] and male singular 3rd person verb can be used for all nouns [53]. Sinhala is also a head-final language, where the complements and modifiers would appear before their heads [55] this is similar to that of English and dissimilar to that of French. In total, according to Abhayasinghe [56], there are 25 types of simple sentence structures in Sinhala. Similar to many Indo Aryan languages, animacy plays a major role in Sinhala grammar in syntactic and semantic roles [57, 58, 59]. Comparative studies done by Noguchi [60] and by Miyagishi [61, 50] have found that animacy extends its influence from phrase level to sentence level in Sinhala (e.g., Usage of post-positions [51, 62]). On this matter, Table I explains grammatical cases and inflections of animate common nouns while Table II explains grammatical cases and inflections of inanimate common nouns. We provide a comparative analysis of parsing the very simple English sentence “I eat a red apple” and its Sinhala, Hindi, and French translations in Fig 4. English and French parsing was done using the Stanford Parser [^7] http://nlp.stanford.edu:8080/parser/. Hindi parsing was done using the IIIT-Hyderabad Parser [^8] http://ltrc.iiit.ac.in/analyzer/ and the study by Singh et al. [63].
(a) English
(b) Sinhala
(c) Hindi
(d) French
@@ 3.1 Corpora
A word corpus of approximately 35,000 entries was developed by Weerasinghe et al. [97]. But it does not seem to be online anymore. Guzmán et al. [98] provided two monolingual corpora for Sinhala. Those were a 155k+ sentences of filtered Sinhala Wikipedia [^14] http://bit.ly/2EQZ7oM and 5178k+ sentences of Sinhala common crawl [^15] http://bit.ly/2ZaQFZo.  Wijeratne and de Silva [99] have publicly released [^16] https://bit.ly/2GEI4d6 a massive corpus of text and stop words taken from a decade of Sinhala Facebook posts. While the stop word extraction of Wijeratne and de Silva [99] algorithmic, Lakmal et al. [100] has introduced a manually curated stop word list [^17] http://bit.ly/3wFi0Wf. A parallel corpus of Sinhala and English was collected by Bañón et al. [101] containing 217,407 sentences and available to download from their website [^18] https://www.paracrawl.eu/. However, the later audit by Caswell et al. [102] raised issues on the quality of that data set. A parallel corpus [^19] https://github.com/kdissa/comparable-corpus of aligned Sinhala-English documents and sentences obtained from crawling the web was released by Sachintha et al. [103]. The study by Warusawithana et al. [104] created a refined version [^20] https://bit.ly/3JF77t9 of the OpenSLR-52 speech corpus for Sinhala [^21] https://openslr.org/52 by Kjartansson et al. [105] [^22] A later further overview on the corpus is also available [106].Dhananjaya et al. [107] created sin-cc-15M corpus2323233https://tinyurl.com/42un7a9y, which they claim, at the time of their publication to be the largest monolingual Sinhala corpus. A text corpus collected from the Sihala blog Kalaya [^24] http://www1.kalaya.org/ along with the relevant code is available on Github [^25] https://bit.ly/KalayaCorpus. A Sinhala NEWS corpus of more than 500,000 articles collected from various online free news websites was released [^26] https://github.com/Sinhala-NLP/NSINA by Hettiarachchi et al. [108]. ImaniGooghari et al. [109] published Glot500272727https://github.com/cisnlp/Glot500, a corpus of 500 languages that includes Sinhala. The have 7,293,178 sentences in their Sinhala corpus. Further, they train a Sinhala language model and report the results of various tasks such as: sentence retrieval, NER, PoS, Text Classification, and round-trip alignment. They also report their result on perplexity analysis.

### v26

@@ Abstract
Sinhala is the native language of the Sinhalese people who make up the largest ethnic group of Sri Lanka. The language belongs to the globe-spanning language tree, Indo-European. However, due to poverty in both linguistic and economic capital, Sinhala, in the perspective of Natural Language Processing tools and research, remains a resource-poor language which has neither the economic drive its cousin English has nor the sheer push of the law of numbers a language such as Chinese has. A number of research groups from Sri Lanka have noticed this dearth and the resultant dire need for proper tools and research for Sinhala natural language processing. However, due to various reasons, these attempts seem to lack coordination and awareness of each other. The objective of this paper is to fill that gap of a comprehensive literature survey of the publicly available Sinhala natural language tools and research so that the researchers working in this field can better utilize the contributions of their peers. As such, we shall be uploading this paper to arXiv and perpetually update it periodically as a living research article to reflect the advances made in the field. This manuscript is at version $5.7.1$.
@@
Sinhala, Natural Language Processing, Low-Resource Languages
@@ 1 Introduction
Sinhala [^1] Englebretson and Genetti [1] observe that in some contexts the Sinhala language is also referred as Sinhalese, Singhala, and Singhalese language, being the native language of the Sinhalese people [2, 3, 4], who make up the largest ethnic group of the island country of Sri Lanka, enjoys being reported as the mother tongue (L1) of approximately 17 million people [5, 6]. When both L1 and L2 speakers are counted, $79.7\%$ of the total Sri Lankan population are literate in Sinhala [7]. A strong correlation between the island of Sri Lanka and the usage of Sinhala has been observed [8] [^2] Dunn and Edwards-Brown [8] in their work to augment Language Identification (LID) task with geography observe that for Sinhala LID does not benefit from the geography model given that the to models already has an alignment of 99%.. This implies two things: 1) the majority of the Sinhala linguistic sources are located in the geographical area of Sri Lanka, and 2) the majority of the available linguistic sources in Sri Lanka are in Sinhala. Sri Lanka is the only country where Sinhala is recognised as an official language [9].
@@
To give a brief linguistic background for the purpose of aligning the Sinhala language with the baseline of English, primarily it should be noted that the Sinhala language belongs to the same Indo-European language tree [10, 11, 12]. However, unlike English, which is part of the Germanic branch, Sinhala belongs to the Indo-Aryan branch. An extremely simplified version of the Indo-European language tree focusing mainly on Sinhala and its close relatives is shown in Fig 1. Further, Sinhala, unlike English, which borrowed the Latin alphabet, has its own writing system, which is a descendant of the Indian Brahmi script [13, 14, 15, 16, 17, 18, 19]. By extension, this makes Sinhala Script a member of the Aramaic family of scripts [20, 21]. Thus by inheritance, the Sinhala writing system is abugida (alphasyllabary), which to say, that consonant-vowel sequences are written as single units [22]. An extremely simplified evolution tree of the Sinhala script is shown in Fig 2. In it, the Sinhala script is shown as a singular derivation from Brahmi directly appearing with the contemporary script. But this is a simplification. We show the actual evolutionary eras of the Sinhala script in Appendix A where the gradual change from Brahmi to the contemporary script is clearer. It should be noted that, due to various historical reasons [23], the modern Sinhala language has loanwords from languages such as Tamil [24], English, Portuguese, and Dutch. An analysis of the distribution of Sinhala letters towards the beginning, middle, or end of words conducted in the style of proposed by ProoffReader [25] using the SinMin Sinhala corpus created by Upeksha et al. [26, 27] is given in Appendix B.
@@
Figure 1: A portion of the Indo-European language tree with a focus on Sinhala and its close relatives. Only the Elu sub-tree is shown exhaustively. All other sub-trees only list some languages for representation purposes, chosen randomly or otherwise. Some of the languages shown as Extinct have archaeological proof (eg: Prakrit) while others were retroactively reconstructed by linguists by observing the commonalities of languages that they begot (eg: Proto-Indo-Aryan). The status of some languages have spread beyond linguistics to be a social and political question. In this depiction, we have attempted to side with academic consensus when available and the majority when not. As an example of one contentions point on this diagram, note the positioning of Pali. While what we show here is the academically accepted relationship, there are other schools of thought: 1) Some argue that Pali is a constructed language (ConLang) of Sinhala and thus should be depicted as a child of Sinhala., 2) Others claim it to be a sibling of Sinhala, 3) Further, some claim it to be either be the link between Elu and Sinhala or the parent of Elu, 4) Some claim it needs to be merged with Gandhari, and 5) Some claim it to be the same as Magadhi Prakrit given that in some Buddhist traditions, it is noted Buddha to have spoken Magadhi.
@@
Figure 2: A portion of the writing systems (scripts) tree with a focus on Sinhala and its close relatives. It is important to note that this is distinct from the language tree shown in Fig 1. The writing systems used by a language being related, not being related, or the degree to which they are related, is not directly correlated to the said languages being related, not being related, or the degree to which they are related. As an example, note that the writing systems of Sinhala and Tamil are reasonably closely related. But Tamil is not shown on Fig 1 given that it does not belong to the Indo-European language family. It belongs to the Dravidian language family. A language that is shown as a single entity in Fig 1 may have multiple entries in this diagram. An example is the language shown as Hindustani in Fig 1 splits into Hindi and Urdu in this diagram, where the former uses the Devanagari script and the latter a derivation of Arabic script. Conversely, languages that may be shown as distinct entities in Fig 1, may map to a singular entity in the script tree. An example is English which would go under Proto-Germanic and Italian which would go under Proto-Italic in an extended version of Fig 1, would both map to Latin in this diagram, given that both of them use the Latin script.
@@
Natural Language Processing (NLP) is a broad area covering all computational processing and analysis of human languages. To achieve this end, NLP systems operate at different levels [37, 38, 39]. A graphical representation of NLP layers and application domains are shown in Figure 3. On one hand, according to Liddy [38], these systems can be categorized into the following layers; phonological, morphological, lexical, syntactic, semantic, discourse, and pragmatic. The phonological layer deals with the interpretation of language sounds. As such, it consists of mainly speech-to-text and text-to-speech systems. In cases where one is working with the written text of the language rather than speech, it is possible to replace this layer with tools which handle Optical Character Recognition (OCR) and language rendering standards (such as Unicode [40]). The morphological layer analyses words at their smallest units of meaning. As such, analysis of word lemmas and prefix-suffix-based inflection is handled in this layer. Lexical layer handles individual words. Therefore, tasks such as Part of Speech (PoS) tagging happens here. The next layer, syntactic, takes place at the phrase and sentence level, where grammatical structures are utilised to obtain meaning. Semantic layer attempts to derive the meanings from the word level to the sentence level. Starting with Named Entity Recognition (NER) at the word level and working its way up by identifying the contexts they are set in until arriving at the overall meaning. The discourse layer handles meaning in textual units larger than a sentence. In this, the function of a particular sentence maybe contextualised within the document it is set in. Finally, the pragmatic layer handles contexts read into contents without having to be explicitly mentioned [37, 38]. Some forms of anaphora (coreference) resolution [41, 42, 43, 44, 45] fall into this application.
@@
The objective of this paper is to serve as a comprehensive survey on the state of natural language processing resources for the Sinhala language. The initial structure and content of this survey are heavily influenced by the preliminary surveys carried out by de Silva [30] and Wijeratne et al. [37]. However, our hope is to host this survey at arXiv as a perpetually evolving, living research article [47, 48] which continuously gets updated as new research and tools for the Sinhala language are created and made publicly available. We also discuss how the non-compliance of policies to put data and code online [49], after the research is concluded and the paper is published, has negatively impacted the growth and sustainability of Sinhala NLP. Hence, it is our hope that this work will help future researchers who are engaged in Sinhala NLP research to conduct their literature surveys efficiently and comprehensively. For the success of this survey, we shall also consider the Sri Lankan NLP tools repository, lknlp [^4] https://github.com/lknlp/lknlp.github.io. This manuscript is at version $5.7.1$. The latest version of the manuscript can be obtained from arXiv [^5] https://arxiv.org/abs/1906.02358 or ResearchGate [^6] http://bit.ly/31AhvvR. Given that this work was initially published in 2019, when citing this work, it is recommended to cite the most recent version to avoid the reviewers questioning its continued contemporary relevance. An example of an updated reference is given in Code Block LABEL:lst:ref.
@article{de2026survey,
@@
journal={arXiv preprint arXiv:1906.02358v26},
year={2026}
@@
The remainder of this survey is organized as follows: Section 2 introduces some important properties and conventions of the Sinhala language, which are important for the development and understanding of Sinhala NLP. Section 3 discusses the various tools and research available for Sinhala NLP. In this section, we would discuss both pure Sinhala NLP tools and research as well as hybrid Sinhala-English work. We will also discuss research and tools that contribute to Sinhala NLP, either along with or with the help of Tamil, the other official language of Sri Lanka. Section 4 gives a brief introduction to the primary language sources used by the studies discussed in this work. Finally, Section 5, concludes the survey.
@@ 2 Properties of the Sinhala Language
Before moving on to discussing Sinhala NLP resources, we shall give a brief introduction to some of the important properties of Sinhala language, which impact the development of Sinhala NLP resources. Sinhala grammar has two forms: written (literary) and spoken. These forms differ from each-other in their core grammatical structures [32, 51, 1, 52]. The written form strictly adheres to the SOV (Subject, Object, and Verb) configuration [53, 54]. Further, in the written form, subject-verb agreement is enforced [55] such that, in order to be grammatically correct, the subject and the verb must agree in terms of: gender (male/female), number (singular/plural) and person (1st/2nd/3rd). However, in spoken Sinhala, the SOV order can be neglected [56] and male singular 3rd person verb can be used for all nouns [55]. Sinhala is also a head-final language, where the complements and modifiers would appear before their heads [57] this is similar to that of English and dissimilar to that of French. In total, according to Abhayasinghe [58], there are 25 types of simple sentence structures in Sinhala and according to Madushani [59] there are 8 types of Verb Phrases. Similar to many Indo Aryan languages, animacy plays a major role in Sinhala grammar in syntactic and semantic roles [60, 61, 62]. Comparative studies done by Noguchi [63] and by Miyagishi [64, 52] have found that animacy extends its influence from phrase level to sentence level in Sinhala (e.g., Usage of post-positions [53, 65]). On this matter, Table I explains grammatical cases and inflections of animate common nouns while Table II explains grammatical cases and inflections of inanimate common nouns. We provide a comparative analysis of parsing the very simple English sentence “I eat a red apple” and its Sinhala, Hindi, and French translations in Fig 4. English and French parsing was done using the Stanford Parser [^7] http://nlp.stanford.edu:8080/parser/. Hindi parsing was done using the IIIT-Hyderabad Parser [^8] http://ltrc.iiit.ac.in/analyzer/ and the study by Singh et al. [66].
((a)) English
((b)) Sinhala
((c)) Hindi
((d)) French
@@ 3.1 Corpora
A word corpus of approximately 35,000 entries was developed by Weerasinghe et al. [100]. But it does not seem to be online anymore. Guzmán et al. [101] provided two monolingual corpora for Sinhala. Those were a 155k+ sentences of filtered Sinhala Wikipedia [^14] http://bit.ly/2EQZ7oM and 5178k+ sentences of Sinhala common crawl [^15] http://bit.ly/2ZaQFZo.  Wijeratne and de Silva [102] have publicly released [^16] https://bit.ly/2GEI4d6 a massive corpus of text and stop words taken from a decade of Sinhala Facebook posts. While the stop word extraction of Wijeratne and de Silva [102] algorithmic, Lakmal et al. [103] has introduced a manually curated stop word list [^17] http://bit.ly/3wFi0Wf. A parallel corpus of Sinhala and English was collected by Bañón et al. [104] containing 217,407 sentences and available to download from their website [^18] https://www.paracrawl.eu/. However, the later audit by Caswell et al. [105] raised issues on the quality of that data set. A parallel corpus [^19] https://github.com/kdissa/comparable-corpus of aligned Sinhala-English documents and sentences obtained from crawling the web was released by Sachintha et al. [106]. The study by Warusawithana et al. [107] created a refined version [^20] https://bit.ly/3JF77t9 of the OpenSLR-52 speech corpus for Sinhala [^21] https://openslr.org/52 by Kjartansson et al. [108] [^22] A later further overview on the corpus is also available [109].Dhananjaya et al. [110] created sin-cc-15M corpus2323233https://tinyurl.com/42un7a9y, which they claim, at the time of their publication to be the largest monolingual Sinhala corpus. A text corpus collected from the Sihala blog Kalaya [^24] http://www1.kalaya.org/ along with the relevant code is available on Github [^25] https://bit.ly/KalayaCorpus. A Sinhala NEWS corpus of more than 500,000 articles collected from various online free news websites was released [^26] https://github.com/Sinhala-NLP/NSINA by Hettiarachchi et al. [111]. Imani et al. [112] published Glot500272727https://github.com/cisnlp/Glot500, a corpus of 500 languages that includes Sinhala. The have 7,293,178 sentences in their Sinhala corpus. Further, they train a Sinhala language model and report the results of various tasks such as: sentence retrieval, NER, PoS, Text Classification, and round-trip alignment. They also report their result on perplexity analysis.